OpenStudy (anonymous):

Find the limit as x approaches infinity of the function: [(2x-3)/(2x+3)]^{2x+1} I did this correctly before, but I suddenly forgot it. I would like some help here.

OpenStudy (anonymous):

probably best bet is to take the log and compute the limit of that

OpenStudy (anonymous):

So it is the fraction raised to the power (2x + 1)

OpenStudy (anonymous):

Ok this is what I did: Let y = that fcn Then, ln(y) = ln(that fcn) Now, I do not know what to do next.

OpenStudy (anonymous):

pretty much when yo see the variable in the exponent, first step is to take the log

OpenStudy (anonymous):

Oops, then also: limit as x-> infinity ln(y) = limit as x-> infinity ln(that fcn)

OpenStudy (anonymous):

Now I forgot what to do next.

OpenStudy (anonymous):

\[(\frac{2x-3}{2x+3})^{2x+1} \] take the log get \[(2x+1)\ln(\frac{2x-3}{2x+3})\]

OpenStudy (anonymous):

now maybe expand or use l'hopital at this step

OpenStudy (anonymous):

How would L'H even work? Wouldn't the inside of ln yield 2?

OpenStudy (anonymous):

Sorry 0.

OpenStudy (anonymous):

OHHHH

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

So you can move the limit into the ln argument right, because ln is continuous for x -> infinity?

OpenStudy (anonymous):

And then you get ln(1), which simplifies to 0.

OpenStudy (anonymous):

And then you move 1/(2x + 1) to the denominator.

OpenStudy (anonymous):

Take the limit of that, which is also 0.

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

you have \[\infty\times 0\] as your form here, so you can write \[(2x+1)\ln(\frac{2x-3}{2x+3})=\frac{\ln\frac{2x-3}{2x+3})}{\frac{1}{2x+1}}\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

oh i am mistaken

OpenStudy (anonymous):

very mistaken this is a lot of work

OpenStudy (anonymous):

maybe this would be easier \[(2x+1)\left(\ln(2x-3)-\ln(2x+3)\right)\]

OpenStudy (anonymous):

no, that is not easier

OpenStudy (anonymous):

1/e^6

OpenStudy (anonymous):

I guess I was wrong ....mathg8 is correct according to wolfram alpha

OpenStudy (anonymous):

I got it !

OpenStudy (anonymous):